It is intended to prepare 2 L of a 0.170 M NaOH solution. For this, NaOH is available 54. 8% ± 0. 3% (m/m), with relative density 1.52 ± 0.01. The pipette used has a ±0.10ml error. The error of the volumetric balloon is 0.005 L. Determine the volume a pipette. Establish the formula that allows finding this value. Invert the formula to obtain the final concentration as a function of the remaining variables. Estimate the error associated with

predicted final concentration from the previous formula


1
Expert's answer
2021-10-11T06:10:19-0400

Molar mass of NaOH = 39.997 g/mol

= 0.17×39.997= 6.8g

2000/6.8= 294.12g

0.548×294.12

= 161.18g